OpenStudy (anonymous):

Factor: 2ax + 3ay + 2bx + 3by Which of the following is one of the factors? (a - b) (2x + 3y) (2x - b)

OpenStudy (anonymous):

Did you factor the equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know thr answer to it

OpenStudy (anonymous):

Do you know how to?

OpenStudy (anonymous):

You can factor it by taking out the GCF which each of them seem to share with each other: \[=a(2x+3y)+b(2x+3y)\]\[=(a+b)(2x+3y)\]

OpenStudy (anonymous):

And thus the second option...

OpenStudy (cwrw238):

= a(2x + 3y ) + b(2x + 3y) so 2x + 3y is a factor